OpenStudy (anonymous):

A town's public works department has ordered sand to spread on roadways during the upcoming winter snows. The delivered sand forms a right cone 36 feet high and a base diameter of 54 feet. How many cubic yards of sand has the department received?

OpenStudy (anonymous):

the formula for the volume of a cone V=(1/3)pi r^2 h, where r=radius and h= height, D=diameter=2r,or r=D/2 V=(1/3)pi(D/2)^2 (36)=8748 pi=27,482.65 cubic feet

OpenStudy (anonymous):

hmm answer wasn't accepted